OpenStudy (anonymous):

Which of the following is a polynomial with roots 4, -2i, and 2i ? Choose one answer. a. x3 - 4x2 + 4x - 16 b. x3 + 4x2 + 4x + 16 c. x3 + 2x2 + 4x + 8 d. x3 - 2x2 + 4x - 8

OpenStudy (anonymous):

Thanks!

OpenStudy (anonymous):

well x^3+4x^2+4x-16 By grouping (x^3+4x^2)+(4x-16) Therefore: x^2(x+4)+4(x-4)=0 so:( X^2+4)(X^2-4)=0 X^2-16=0 so. X^2=16 . So X=+or- 2

OpenStudy (anonymous):

& if u tried this method on the other 3 equations i think the answer will be C